The measure of a supplement of <1 is six times the measure of a complement of <1. Find the measure of <1, its supplement, and its complement.

Answer by mananth(16946)

let angle be x
supplement = 180-x
complement = 90-x
...
180-x=6(90-x)
180-x=540-6x
-x+6x=540-180
5x=360
x= 360/5
x= 72 deg. the angle